two cars travel the same distance. car 1 travels at 50 mph and car 2 at 65 mph. If it takes car 1 one more hour to travel than car 2 how far did the cars travel? remember d= rt
a. write an expression to represent the distance car 2 travels
b. write an expression to represent the distance car 1 travels
c. use these expressions to find time car 2 traveled
d. use this information to determine the distance traveled

Answer by edjones(8007)

rt=rt because d[1]=d[2]
50(t+1)=65t
50t+50=65t
15t=50
t=10/3=3:20
65 * 10/3 = 216 2/3 mi. Distance they traveled.
.
Ed